OpenStudy (anonymous):

What is the z-value of the solutions to the following system of equations? x+y-z=0 x -2z=-1 2x+2y+z=9

OpenStudy (anonymous):

z=3

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

@brittany1696 2x+2y+z=9 and -2(x+y-z)=0,after we will get 3z=9 then z=3, x-6=-1 so x=5 , so 5+y-3=0 so y=-2